# What Evidence Do Language Models Find Convincing?

### v1

@@ 1 Introduction
We perform sensitivity and counterfactual analyses to find in-the-wild features that correlate with document convincingness. We consider a mix of features that describe (1) stylistic properties of a document and (2) relevance of a document to the question. Many of these were inspired by results from studies of human credibility, for example, we consider whether adding scientific references makes text appear more convincing.
@@ 2 Background and Motivations
Figure 2: Models over-rely on document relevance. We study how the convincingness of a particular evidence paragraph (measured through win-rate) changes when we modify it. We compare the effect of these changes to a baseline perturbation where we append “Thanks for reading!” to the end (indicated by the dotted line). We find that many stylistic changes—inspired by factors that influence humans—have a neutral or even negative effect on models. On the other hand, perturbations that increase the texts relevance but minimally change its style have a substantial positive effect on models. Descriptions for each perturbation can be found in Appendix E.
@@ 3 The ConflictingQA Dataset
Category
Example Question
Num Evidence Docs
Pharmacology
Are antidepressants more effective than placebo?
10
Online Learning
Are online degrees valued less by employers?
10
Biodiversity
Are bees the most important pollinators?
10
Web Design
Does longer website content rank better on Google?
13
Sustainability
Are electric cars really green?
9
Philosophy
Are humans fundamentally good or evil?
7
Nuclear Energy
Can nuclear power solve climate change?
7
Work-Life Balance
Is unlimited vacation time beneficial for employees?
10
Somnology
Do older people need less sleep?
8
Biomechanics
Do compression garments improve athletic performance?
13
Table 1: In ConflictingQA, we create controversial questions for 136 different categories (see Table 5 for the complete list). Above, we show an example question for ten different categories, as well as the number of evidence paragraphs for each one. The evidence paragraphs contain a mix of Yes and No answers.
@@
Collecting contentious questions. We first create a series of realistic open-ended questions for which there exists conflicting evidence online. Critically, unlike past work on ambiguity in QA (Min et al., 2020; Zhang and Choi, 2021; Sun et al., 2023), we want to collect unambiguous questions that still have answer conflicts. For example, in Figure 1, we show a question “are artificial sweeteners linked to cancer?”, which is a widely-debated query in which there exist websites that support both answers. We design the questions to elicit binaries responses of Yes or No to simplify evaluation.
@@
Collecting evidence paragraphs. Given these questions, we want to find evidence paragraphs that support both the answers of Yes and No. We also want these paragraphs to (1) contain a diverse range of argument styles, factual information, etc., and (2) be realistic inputs to an LLM. To handle this, we emulate running an real-world retrieval-augmented LLM system that uses the Google Search API as its retrieval engine. Concretely, we take the user’s query, reformulate it, and take the top-$k$ results from Google search for the answer Yes and the top-$k$ for the answer No.
@@
We first turn each question into affirmative and negative statements, e.g., the question “is asparatame safe?” is converted to “asparatame is safe” and “asparatame is harmful” using GPT-4. We also put double quotes (to indicate to Google Search that we have exact-match keywords) around any tokens that do not change after rephrasing the question into either statements (e.g., “aspartame”). For both the affirmative and negative statements, we search the queries using the Google Search API and retrieve top-$k$ documents. [^1] We set $k=20$ because qualitatively the relevancy of the results dropped off significantly after this point. As is common in many retrieval-augmented models (Nakano et al., 2021), we do not consider any visual features of the web page. Instead, we extract the raw text from each document using jusText. [^2] Package available at https://github.com/miso-belica/jusText. Although humans use visual features when considering the credibility and trustworthiness of a source Kakol et al. (2017); Fogg et al. (2003), we do not consider these features as most state-of-the-art LLMs do not use visual inputs. Additionally, we do not explicitly include metadata like source URL, publication date, or page headings.
When searching queries such as “aspartame is safe”, we still retrieve documents that argue that aspartame is unsafe. To label the documents actual stance, we use an ensemble of claude-instant-v1 and GPT-4-1106-preview and keep only the samples where the two models agree (see Table 7 in Appendix A for the prompts). [^3] After identifying the stance, we also feed the paragraphs into the downstream LLM that we are testing and make sure that its answer aligns with the paragraphs predicted stances. This further filters and balances the data, accounting for mistakes in the downstream model. See Appendix B for details. Furthermore, we allow the LLM to say that a document is irrelevant to the query; if so, we also filter it from the input.
@@ 4.1 Convincingness as Paragraph Win Rate
Models Cannot Predict Convincingness We designed the above experimental setting to emulate how production RAG models work. However, we could have instead just directly asked the LLM, “do you find paragraph X to be persuasive?”. This is how humans are typically asked to judge the convincingness of a piece of evidence (Kakol et al., 2017; Jo et al., 2019; Kakol et al., 2013). However, we find that LLMs are largely incapable of expressing the convincingness of a paragraph in words, e.g., there is little correlation in which paragraphs are marked as convincing in the two settings (Figure 3). [^4] Our methodology for this setting is described in more detail in Appendix D. We thus focus on the more practically-grounded setting going forward.
@@ 4.2 Implementation Details
We evaluate a mix of open-source (LLaMA-2 Chat (Touvron et al., 2023b), Vicuna v1.5 (Chiang et al., 2023), and WizardLM v1.2 (Xu et al., 2023)) and closed source (GPT-4, Anthropic Claude v1 Instant) models. Importantly, we specify “Use only the information in the above text to answer the question” as we are looking to see how models judge stylistic differences in evidence, rather than their prior stances on the question.
We extract binary Yes/No predictions from the model for each question. For open-source models, we compare the log-probabilities of the next-token. For the closed-source models, we prompt them to output only Yes or No. See Table 8 for the prompt used for question answering.
Figure 3: Humans can read a paragraph in isolation and evaluate how convincing it is. For LLMs, when they are given a paragraph in isolation, they are unable to express its convincingness in words. Concretely, we plot the win rate of paragraphs versus what a model outputs when it is asked to judge the convincingness on a 1–5 Likert scale. The error bars show a 95% CI.
@@
Figure 4: Why do models prefer certain paragraphs over others? We test correlations between different features and paragraph win-rates. Here, we show LLaMA-2 Chat 13B (see all other models in Appendix C), where the model tends to prefer samples with low-perplexity 4(d). In addition, paragraphs with high relevancy scores—particularly high question-paragraph embedding similarity are also highly convincing 4(f). See Figure 2 for additional analysis. The error bars show the 95% CI (n = 242).
@@ 4.3 What Correlates With Convincingness?
Results Stylistic features are poor predictors of paragraph convincingness. Figure 4 shows the results for the LLaMA-2 Chat model and Figures 58 shows the results for other models. For example, across all models the Flesch-Kincaid score and number of unique tokens does not correlate with convincingness. Similarly, paragraphs with a more positive sentiment and lower perplexity tend to have some small impact on convincingness, with varying strengths from model to model.
On the other hand, question-paragraph embedding similarity correlates strongly with win-rate across all models except for GPT-4. Similarly, a positive (but weaker) correlational exists between n-gram overlap and win-rate.
@@ 4.4 Counterfactual Analysis
Rather than a correlational study, we also test how win-rates change in a counterfactual setting where we directly edit paragraphs using an LLM. We make perturbations using claude-v1-instant, examples of which are shown in Figure 9.
Stylistic changes We first consider changes inspired by factors that humans find important for the credibility of a text. For example, adding more information, adding scientific references, or making the text sound more objective. Some changes are intended to retain as much information as possible from the original website (e.g., Add Science Reference, Add More Info). Others involve significantly changing the entire paragraph (e.g., Rewrite Objective, Rewrite Tech. Language). All of the perturbations are described further in Appendix E.
Relevancy changes Based on the results in Section 4.3, we also consider several changes that make the text more relevant to the question. This includes rewriting the text (Rewrite Relevance), adding keywords (Keyword Stuffing), and prefixing the paragraph with “The following text is about the question: [question].” (Question Prefix). Finally, we consider a perturbation inspired by the “AddSent” perturbation in Jia and Liang (2017) where we use claude-v1-instant to add a single sentence to make stance of a text obvious (Add Single Sentence). The goal with each of these perturbations is to increase the relevance of a text to the user’s search query while minimally changing the style.
We also compare these perturbations against a “control” perturbation where text is suffixed with “Thanks for reading!” This perturbation minimally influences both style and relevance. For simplicity we only perturb the paragraphs with the Yes stance.
Counterfactual results The results for the counterfactual experiments are shown in Figure 2: compared to the effect of the control perturbation, stylistic features tend to have a neutral to negative effect while relevancy-based features significantly improve win-rate. Note that many of these perturbations change a smaller amount of tokens than stylistic features—leaving the content of the website largely unchanged (e.g., Add Single Sentence, Question Prefix)—but are still able to improve the convincingness of websites.
Overall, we find that, as compared to typical finding from human experiments Fogg et al. (2003); Kakol et al. (2013); Metzger et al. (2010), LLMs tend to overindex on relevancy. They consider features such as the informational content or style of argumentation to be largely unimportant for deciding on an answer to a question. Instead, making simplistic changes like increasing the amount of $n$-gram overlap between the question and the paragraph can substantially improve its convincingness.
@@
Are Coral snakes found in Africa?
Old-world coral snakes are found in Africa, the Middle East, India, and parts of Southeast Asia. New World coral snakes can be found in North America, Central America, and South America.
Coral snakes are found in scattered localities in the southern coastal plains from North Carolina to Louisiana, including all of Florida.
Are Florida Panthers on the brink of extinction?
As Florida’s panther numbers plummeted, the state’s human population nearly doubled over the past 30 years. Recent development patterns pose threats to panthers.
Now, though, their population is on the upswing … Both the numbers and the genetic diversity of Florida panthers improved.
Are artificial sweeteners safe for diabetics?
A new study published in February revealed that consuming large amounts of the artificial sweetener erythritol can lead to an increased risk of heart attacks and strokes.
Furthermore, xylitol does not need insulin to be metabolized, so it can be safely consumed by diabetics.
Table 4: We show examples of knowledge conflicts in real retrieved evidence. For example, questions may be underspecified (e.g., “old-world” vs “new-world” coral snakes). In other cases, the answer is dependent on the publication date (e.g., currently on the brink vs recent upswing). Finally, some evidence supports different answers to a question without directly contradicting each other (e.g., the safety of two different artificial sweeteners).
@@ 5 Discussion & Related Work
How should systems handle ambiguity? One reasonable suggestion is that agents should not make their own autonomous decisions when faced with ambiguous or conflicting evidence. For example, they may summarize both sides of the aspartame argument, or they may ask the user to clarify their preferences. There is naturally a trade-off between autonomy and clarity. Past work has explored one side of this trade-off, for example by abstaining from answering in cases of ambiguity Chen et al. (2022), by trying to provide multiple perspectives on the answer Min et al. (2020), or by asking clarification questions (Rao and Daumé III, 2018; Zamani et al., 2020). Our work explores the other side of the trade-off: we analyze the behavior of models when they are expected to resolve ambiguity with more autonomy.
Additionally, our dataset serves as a benchmark for exploring these questions as it reflects real-world ambiguities in question-answering. For example, in Table 4) to best answer “Are Coral snakes found in Africa?”, additional clarification questions would be needed from the user.
@@
Improving model judgements. Our work highlights the gap between human and model judgements of text credibility. This solution to this, however, is not clear cut. For one, it is not clear the level of discretion models should have when making predictions. Human judgements of website credibility differ from person to person (Kakol et al., 2013), and users may not be comfortable with the idea that models are “choosing” for them what source to trust. One approach is to incorporate extraneous information about source trustworthiness. For example, Bashlovkina et al. (2023) propose aligning model predictions with that of known trustworthy sources via prompting. Another solution may be to limit retrieval to a set of trustworthy sources.
@@ 6 Conclusion
We study how RAG model judge convincingness by collecting a diverse set of controversial questions and website text (ConflictingQA), and designing a realistic evaluation framework based on how these models are used in practice. Our results show that today’s LLMs tend to overrely on relevancy and ignore many stylistic features of text that humans often deem important. Future work should explore how integrating other forms of information (e.g., metadata, visual content) can influence these behaviors. In addition, given the possible flood of LLM-generated content on the internet, it is important to consider how these synthetic texts may influence LLM judgements of convincingness.
@@
While ConflictingQA is diverse and simulates real-world uses of RAG models, it may not fully capture the complexity of how LLMs are used in practice. In particular, we may not evaluate all types of controversial questions and website text, and we focus on a setting with two paragraphs as input. We also only consider a binary Yes or No answer to contentious questions whereas LLM outputs in practice may be more nuanced. Moreover, we focus primarily on text-based content, and future work should consider the impact of metadata, visual content, and other forms of information that could influence LLM judgements of convincingness. Finally, we acknowledge that our study does not address the broader ethical and societal implications of LLMs both reading and generating most of the content on the web. Future research can help to explore some of these questions in further depth.

### v2

@@ 1 Introduction
We perform sensitivity and counterfactual analyses to find in-the-wild features that correlate with document convincingness. We consider a mix of features that describe (1) stylistic properties of a document and (2) the relevance of a document to the question. Many of these were inspired by results from studies on human credibility judgements: for example, we consider whether adding scientific references makes text appear more convincing.
@@ 2 Background and Motivations
Figure 2: Models over-rely on document relevance. We study how the convincingness of a particular evidence paragraph (measured through win-rate) changes when we modify it. We compare the effects of these changes to a baseline perturbation where we append “Thanks for reading!” to the end of the text (indicated by the dotted line). We find that many stylistic changes—inspired by factors that influence humans—have a neutral or even negative effect on models. On the other hand, perturbations that increase the text’s relevance but minimally change its style have a substantial positive effect on models. Descriptions for each perturbation can be found in Appendix E.
@@ 3 The ConflictingQA Dataset
Category Example Question Num Evidence Docs
Pharmacology Are antidepressants more effective than placebo? 10
Online Learning Are online degrees valued less by employers? 10
Biodiversity Are bees the most important pollinators? 10
Web Design Does longer website content rank better on Google? 13
Sustainability Are electric cars really green? 9
Philosophy Are humans fundamentally good or evil? 7
Nuclear Energy Can nuclear power solve climate change? 7
Work-Life Balance Is unlimited vacation time beneficial for employees? 10
Somnology Do older people need less sleep? 8
Biomechanics Do compression garments improve athletic performance? 13
Table 1: In ConflictingQA, we create controversial questions for 191 different categories (see Table 5 in Appendix A for the complete list). Above, we show an example question for ten different categories, as well as the number of evidence paragraphs used for each question when evaluating LLaMA-2 Chat.
@@
Collecting contentious questions. We first create a series of realistic open-ended questions for which there exists conflicting evidence online. Critically, unlike past work on ambiguity in QA (Min et al., 2020; Zhang and Choi, 2021; Sun et al., 2023), we want to collect unambiguous questions that still have answer conflicts. For example, in Figure 1, we show a question “are artificial sweeteners linked to cancer?”, which is a widely-debated query in which there exist websites that support both answers. We design the questions to elicit binary responses of Yes or No to simplify evaluation.
@@
Collecting evidence paragraphs. Given these questions, we want to find evidence paragraphs that support both the answers of Yes and No. We also want these paragraphs to (1) contain a diverse range of argument styles, factual information, etc., and (2) be realistic inputs to an LLM. To handle this, we emulate running an real-world retrieval-augmented LLM system that uses the Google Search API as its retrieval engine. Concretely, we take the user’s query, reformulate it, and take the top-$k$ results from Google search for the answer Yes and the top-$k$ results for the answer No.
@@
We first turn each question into affirmative and negative statements using GPT-3.5 Turbo, e.g., the question “is aspartame safe?” is converted to “Aspartame is safe. Aspartame isn’t harmful.” and “Aspartame is harmful. Aspartame isn’t safe.”. We also put double quotes (to indicate to Google Search that we have exact-match keywords) around any tokens that do not change after rephrasing the question into either statement (e.g., “aspartame”). For both the affirmative and negative statements, [^1] We find that there’s significant overlap between the webpages returned by the affirmative and negative rephrasings of the query. In future iterations, it would be reasonable to perform searches for only the affirmative or negative statements. we search the queries using the Google Search API and retrieve the top-$k$ documents. [^2] We set $k=20$ because qualitatively the relevancy of the results dropped off significantly after this point. As is common in many retrieval-augmented models (Nakano et al., 2021), we do not consider any visual features of the web page. Instead, we extract the raw text from each document using jusText. [^3] Package available at https://github.com/miso-belica/jusText. Although humans use visual features when considering the credibility and trustworthiness of a source Kakol et al. (2017); Fogg et al. (2003), we do not consider these features as most state-of-the-art LLMs do not use visual inputs. Additionally, we do not explicitly include metadata like source URL, publication date, or page headings.
When searching queries such as “aspartame is safe”, we still retrieve documents that argue that aspartame is unsafe. To label the document’s actual stance, we use an ensemble of claude-instant-v1 and GPT-4-1106-preview and keep only the samples where the two models agree (see Table 9 in Appendix A for the prompts). [^4] After identifying the stance, we also feed the paragraphs into the downstream LLM that we are testing and make sure that its answer aligns with the paragraphs’ predicted stances. This further filters the data, accounting for mistakes in the downstream model. See Appendix B for details. Furthermore, we allow the LLM to say that a document is irrelevant to the query; if so, we also filter it from the input.
@@ 4.1 Convincingness as Paragraph Win Rate
Models Cannot Predict Convincingness. We designed the above experimental setting to emulate how production RAG models work. However, we could have instead just directly asked the LLM, “do you find paragraph X to be persuasive?”. This is how humans are typically asked to judge the convincingness of a piece of evidence (Kakol et al., 2017; Jo et al., 2019; Kakol et al., 2013). However, we find that LLMs are largely incapable of expressing the convincingness of a paragraph in words, e.g., there is little correlation in which paragraphs are marked as convincing in the two settings (Figure 3). [^5] Our methodology for this setting is described in more detail in Appendix D. We thus focus on the more practically-grounded setting going forward.
@@ 4.2 Implementation Details
We evaluate a mix of open-source (LLaMA-2 Chat (Touvron et al., 2023b), Vicuna v1.5 (Chiang et al., 2023), and WizardLM v1.2 (Xu et al., 2023)) and closed source (GPT-4, Anthropic Claude v1 Instant) models, extracting binary Yes/No predictions for each question. Importantly, we specify “Use only the information in the above text to answer the question” as we are looking to see how models judge stylistic differences in evidence, rather than for their prior stances on the question. For open-source models, we compare the log-probabilities of the next-token. For the closed-source models, we prompt them to output only Yes or No. See Table 10 in Appendix A for the prompt used for question answering.
When conducting sensitivity analyses, we also balance the dataset such that each answer (Yes or No) to a question contains an equal number of convincing and unconvincing paragraphs. This way, if the model has a systematic bias toward a certain answer, it would affect the win-rate for convincing and unconvincing paragraphs equally.
Figure 3: Humans can read a paragraph in isolation and evaluate how convincing it is. For LLMs, when they are given a paragraph in isolation, they are unable to express its convincingness in words. Concretely, we plot the win rate of paragraphs versus what a model outputs when it is asked to judge the convincingness on a 1–5 Likert scale, the x-axis representing the quantiles of this metric. The error bars show a 95% CI.
@@
Figure 4: Why do models prefer certain paragraphs over others? We test correlations between different features and paragraph win-rates. Here, we show LLaMA-2 Chat 13B (see all other models in Appendix C), where the model tends to have a slight preference toward samples with low-perplexity 4(d). In addition, paragraphs with high relevancy scores (high question-paragraph embedding similarity) are significantly more convincing 4(f). See Figure 2 for additional analysis. The error bars show the 95% CI (n = 242), and the x-axes represent the quantiles of the target feature.
@@ 4.3 What Correlates With Convincingness?
Results. Stylistic features are poor predictors of paragraph convincingness. Figure 4 shows the results for the LLaMA-2 Chat model and Figures 58 in Appendix C show the results for other models. Across all models, the Flesch-Kincaid score and the number of unique tokens does not correlate with convincingness. Perplexity and sentiment tend to have some small impact on convincingness, with varying strengths from model to model. On the other hand, question-paragraph embedding similarity correlates strongly with win-rate across all models except for GPT-4, and a positive (but weaker) correlation exists between n-gram overlap and win-rate.
@@ 4.4 Counterfactual Analysis
In addition to a correlational study, we also test how win-rates change in a counterfactual setting where we directly edit paragraphs using an LLM. We make perturbations using claude-v1-instant, examples of which are shown in Figure 9 in Appendix E.
Stylistic changes.
We first consider changes inspired by factors that humans find important for the credibility of text by, for example, adding more information, adding scientific references, or making the text sound more objective. Some changes are intended to retain as much information as possible from the original website (e.g., Add More Info). Others involve significantly changing the entire paragraph (e.g., Rewrite Objective, Rewrite Tech. Language). All of the perturbations are described further in Appendix E.
Relevancy changes.
Based on the results in Section 4.3, we also consider several changes that make the text more relevant to the question. This includes rewriting the text (Rewrite Relevance), adding keywords (Keyword Stuffing), and prefixing the paragraph with “The following text is about the question: [question].” (Question Prefix). Finally, we consider a perturbation inspired by the “AddSent” setting in Jia and Liang (2017) where we use claude-v1-instant to add a single sentence to make the stance of a text obvious (Add Single Sentence). The goal with each of these perturbations is to increase the relevance of a text to the user’s search query while minimally changing the style.
We additionally compare these against a “control” perturbation where text is suffixed with “Thanks for reading!” This perturbation minimally influences both style and relevance. For simplicity we only perturb the paragraphs with the Yes stance.
Counterfactual results. The results for the counterfactual experiments are shown in Figure 2: compared to the effect of the control perturbation, stylistic features tend to have a neutral to negative effect while relevancy-based features significantly improve win-rate. Note that many of these perturbations change a smaller amount of tokens than stylistic features—leaving the content of the website largely unchanged (e.g., Add Single Sentence, Question Prefix)—but are still able to improve the convincingness of websites.
Overall, we find that, as compared to typical results from human experiments Fogg et al. (2003); Kakol et al. (2013); Metzger et al. (2010), LLMs tend to overindex on relevancy. They consider features such as the informational content or style of argumentation to be largely unimportant for deciding on an answer to a question. Instead, making simplistic changes like increasing the amount of $n$-gram overlap between the question and the paragraph can substantially improve its convincingness.
@@
Are Coral snakes found in Africa? Old-world coral snakes are found in Africa, the Middle East, India, and parts of Southeast Asia. New World coral snakes can be found in North America, Central America, and South America. Coral snakes are found in scattered localities in the southern coastal plains from North Carolina to Louisiana, including all of Florida.
Are Florida Panthers on the brink of extinction? As Florida’s panther numbers plummeted, the state’s human population nearly doubled over the past 30 years. Recent development patterns pose threats to panthers. Now, though, their population is on the upswing … Both the numbers and the genetic diversity of Florida panthers improved.
Are artificial sweeteners safe for diabetics? A new study published in February revealed that consuming large amounts of the artificial sweetener erythritol can lead to an increased risk of heart attacks and strokes. Furthermore, xylitol does not need insulin to be metabolized, so it can be safely consumed by diabetics.
Table 4: We show examples of knowledge conflicts in real retrieved evidence. For example, questions may be underspecified (e.g., “old-world” vs “new-world” coral snakes). In other cases, the answer is dependent on the publication date (e.g., currently on the brink vs recent upswing). Finally, some evidence support different answers to a question without directly contradicting each other (e.g., the safety of two different artificial sweeteners).
@@ 5 Discussion & Related Work
How should systems handle ambiguity?
One reasonable suggestion is that agents should not make their own autonomous decisions when faced with ambiguous or conflicting evidence. For example, they may summarize both sides of the aspartame argument, or they may ask the user to clarify their preferences. There is naturally a trade-off between autonomy and clarity. Past work has explored one side of this trade-off, for example by abstaining from answering in cases of ambiguity Chen et al. (2022), by trying to provide multiple perspectives on the answer Min et al. (2020), or by asking clarification questions (Rao and Daumé III, 2018; Zamani et al., 2020). Our work explores the other side of the trade-off: we analyze the behavior of models when they are expected to resolve ambiguity with more autonomy.
Additionally, our dataset serves as a benchmark for exploring these questions as it reflects real-world ambiguities in question-answering. For example, in Table 4, to best answer “Are Coral snakes found in Africa?”, additional clarification questions would be needed from the user.
@@
Improving model judgements.
Our work highlights the gap between human and model judgements of text credibility. The solution to this, however, is not clear cut. For one, it is not clear the level of discretion models should have when making predictions. Human judgements of website credibility differ from person to person (Kakol et al., 2013), and users may not be comfortable with the idea that models are “choosing” for them what source to trust. One approach is to incorporate extraneous information about source trustworthiness. For example, Bashlovkina et al. (2023) propose aligning model predictions with that of known trustworthy sources via prompting. Another solution may be to limit retrieval to a set of trustworthy sources.
@@ 6 Conclusion
We study how RAG models judge convincingness by collecting a diverse set of controversial questions and website text (ConflictingQA), and designing a realistic evaluation framework based on how these models are used in practice. Our results show that today’s LLMs tend to overrely on relevancy and ignore many stylistic features of text that humans often deem important. Future work should explore how integrating other forms of information (e.g., metadata, visual content) can influence these behaviors. In addition, given the possible flood of LLM-generated content on the internet, it is important to consider how these synthetic texts may influence LLM judgements of convincingness.
@@
While ConflictingQA is diverse and simulates real-world uses of RAG models, it may not fully capture the complexity of how LLMs are used in practice. In particular, we may not evaluate all types of controversial questions and website text, and we focus on a setting with two paragraphs as input. We also only consider a binary Yes or No answer to contentious questions whereas LLM outputs in practice may be more nuanced. Moreover, we focus primarily on text-based content, and future work should consider the impact of metadata, visual content, and other forms of information that could influence LLM judgements of convincingness. We also acknowledge that our study does not address the broader ethical and societal implications of LLMs both reading and generating most of the content on the web. Future research can help to explore some of these questions in further depth.
Finally, it’s important to note that the study of model robustness is naturally dual-use as adversaries can misuse these insights to exploit RAG systems (for example, by creating “trustworthy” websites containing misinformation). However, as these vulnerabilities can already be exploited in existing models, we instead believe that it’s best for these issues to known and understood by the broader scientific community. Furthermore, by releasing ConflictingQA, we can work toward fixing these vulnerabilities and improving the quality of information generated by RAG systems in general.